the average local cell phone call length was reported to be 2.27 minutes. A random sample of 20 phone calls showed an average of 2.98 minutes with a standard deviation of 0.98 minute. At 0.05 level of significance can it be concluded that the average differs from the population average?
Answer by stanbon(75887)

the average local cell phone call length was reported to be 2.27 minutes. A random sample of 20 phone calls showed an average of 2.98 minutes with a standard deviation of 0.98 minute. At 0.05 level of significance can it be concluded that the average differs from the population average?
-----
Ho: p = 2.27 (claim)
Ha: p # 2.27
--------------------
x-bar = 2.98
----
t(2.98) = (2.98-2.27)/[0.98/sqrt(20)] = 3.24
------
p-value = 2*P(t > 3.24 when df = 19) = 2*tcdf(3.24,100) = 0.0051
---------
Conclusion: Since the p-value is less tahn 5% reject Ho.
The test results do not support the claim at the 5%
level of significance.
========
Cheers,
Stan H.
================